A 12 liter mixture is 60% alcohol. If 8 liters of a 25% alcohol mixture are added,what percent mixture will be formed?

Answer by mananth(16946)

Let x be the percent age of the mixyure
Percent ---------------- quantity
Alcohol I 60 ---------------- 12 liters
Alcohol II 25 ---------------- 8 liters
Total x ---------------- 20 liters

15*60+8 *25=20x
900+200 =20x
1100=20x
/ 20
1100 / 20 = x
55 = x
the percent age of the mixyure is 55